### v2

@@ Abstract
In recent years, consumer Internet of Things (IoT) devices have become widely used in daily life. With the popularity of devices, related security and privacy risks arise at the same time as they collect user-related data and transmit it to various service providers. Although China accounts for a larger share of the consumer IoT industry, current analyses on consumer IoT device traffic primarily focus on regions such as Europe, the United States, and Australia. Research on China, however, is currently relatively rare. This study constructs the first large-scale dataset about consumer IoT device traffic in China. Specifically, we propose a fine-grained traffic collection guidance covering the entire lifecycle of consumer IoT devices, gathering traffic from 77 devices spanning 38 brands and 12 device categories. Based on this dataset, we analyze traffic destinations and encryption practices across different device types during the entire lifecycle and compare the findings with the results of other regions. Compared to other regions, our results show that consumer IoT devices in China rely more on domestic services and overall perform better in terms of encryption practices. However, there are still 23/40 devices improperly conducting certificate validation, and 2/70 devices use insecure encryption protocols. To facilitate future research, we open-source our traffic collection guidance and make our dataset publicly available.
@@
Chinese Consumer IoT, Traffic Dataset, Traffic Measurement, Privacy, Security
@@ 1 Introduction
The consumer Internet of Things (IoT) industry has rapidly developed in recent years, with various devices becoming widespread and offering numerous conveniences. However, the rapid spread of devices has also raised growing concerns regarding security and privacy risks. The data collected by these devices is often transmitted to manufacturers of devices and other service providers. At the same time, data transmission is facing strict legal regulations [1] in China. Therefore, it is necessary to measure consumer IoT traffic from security and privacy perspectives.
@@
Researchers have conducted studies on consumer IoT device traffic in various regions, including Europe [3, 4], the United States [5, 6, 7], and Australia [8, 9], and have identified several security and privacy risks associated with these devices. However, according to IDC [2], China’s consumer IoT industry is projected to reach \$300 billion by 2027. Despite this rapid growth, research on consumer IoT device traffic in China remains limited and lacks comprehensive analysis. This research gap drives us to build a large-scale and comprehensive dataset of consumer IoT traffic in China and conduct a measurement of the dataset.
@@
While constructing the dataset, we introduce methodological improvements compared to existing studies. Previous studies have primarily focused on the device interaction phase [5, 13]. However, we find differences in traffic across different operational phases of the device, and limiting the focus on the interaction phase may lead to an incomplete analysis. Therefore, we propose a fine-grained traffic collection guidance that covers the entire lifecycle of the device to address this limitation. The entire lifecycle refers to all phases during normal device usage, including device setup, user interaction, idle, and deletion [10]. Compared with prior works, the traffic we collected shows greater coverage and can support more fine-grained analysis. Furthermore, we open-source our traffic collection guidance and make our dataset publicly available to facilitate further research.
@@
According to our traffic collection guidance, we collected the first part of device traffic between March 1 and June 1, 2024, covering 50 consumer IoT devices. To further expand the dataset, we conducted two additional rounds of traffic collection: the second from November 1 to November 29, 2024, and the third from June 7 to June 19, 2025, involving 27 more devices. During the traffic collection experiment, which lasted approximately three months in total, we built a dataset comprising 108 hours of traffic from Chinese consumer IoT devices. The total size of the dataset is 3 GB. Compared with the existing datasets, our dataset covers the full lifecycle traffic of the devices and contains the largest number of devices (within a single region) in the dataset. Furthermore, for devices that can update firmware during the experiment, we also collected traffic after their firmware update to study the differences in device traffic before and after the firmware update. Based on the traffic collected, we analyzed the traffic destinations and encryption practices and then compared our findings with those of the United States and the United Kingdom [5, 13].
@@
Our study reveals that, regarding traffic destinations, Chinese devices show a higher reliance on domestic services. The distribution of traffic destinations for Chinese devices is mainly directed to China (99.75%) and relies on domestic organizations. In contrast, devices from the United Kingdom have a significantly higher proportion of overseas traffic, with a greater reliance on organizations from the United States. Regarding encryption practices, devices from different regions all show good encryption practices, although some issues remain. In particular, Chinese devices have less PII exposure in unencrypted traffic and a lower percentage of unencrypted traffic compared to devices in the United States and the United Kingdom. However, 23/40 devices still use improper certificate verification, and 2/70 use insecure encryption protocols.
@@
  • The first Chinese consumer IoT devices traffic dataset. We constructed a dataset that includes 108 hours of traffic from 77 devices across 12 categories and 38 brands. We have publicly released the dataset111[https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to support future research within the community.
@@
  • Security and privacy measurement of Chinese consumer IoT device traffic. We developed scripts to analyze traffic destinations and encryption practices, then compared the results with those from other regions, revealing the characteristics of Chinese consumer IoT devices and regional differences.
@@ 2 Background
Consumer IoT devices typically refer to IoT devices designed for personal or household use, such as cameras, speakers, and lights. These devices can be bound to and controlled by the companion apps, following similar operational phases (i.e., lifecycle). During these phases, users can control the device through various methods. Meanwhile, consumer IoT devices communicate frequently with cloud servers during operation to support various device functions. This section introduces the concepts and definitions relevant to the lifecycle, control methods, and cloud services mentioned above.
@@
The lifecycle of consumer IoT devices. The lifecycle of consumer IoT devices contains four phases based on a prior work [10]: setup, interaction, idle, and deletion. When users first activate a consumer IoT device, the device will enter the setup phase. During this phase, users need to complete operations such as network configuration and bind the device to a companion app. The interaction phase refers to the phase in which users interact with the device, such as controlling light switches and checking camera status. After setting up the device, we refer to the phase in which the user no longer interacts with the device as the idle phase. Finally, the deletion phase refers to the phase when users unbind the device from the companion app, restoring it to its factory settings.
@@
Control methods of consumer IoT device. Common control methods of consumer IoT devices include four types: Wide Area Network (WAN) control, Local Area Network (LAN) control, physical control, and multimodal control. WAN control refers to scenarios where the controlling smartphone with the companion app and the consumer IoT device are not on the same local network. In this case, commands between the app and the device must be relayed through a cloud server. For example, when a user is away from home during travel, he/she can monitor indoor environments in real-time via a camera through a companion app. LAN control occurs when the controlling smartphone and the consumer IoT device are on the same local network. In this case, communications between the device and the app can be established through local protocols, such as Bluetooth. For example, when a user is at home, he/she can adjust the brightness and color temperature of smart lights through a companion app. Physical control involves direct interaction with the device through physical buttons. For example, users can press physical switch buttons on the device to reset or power it off. Consumer IoT devices with various sensors support multimodal control through different methods. Examples include users waking up speakers through voice commands or interacting with devices via triggers from motion sensors.
@@
Cloud servers of consumer IoT devices. When a consumer IoT device operates, the data it collects is shared with various cloud servers. Following the classification approach used in the study by Ren et al. [5], we categorize these servers into three groups: first party, support party, and third party. First party refers to the manufacturer of the consumer IoT device or the developer of its companion app. Support party refers to organizations that provide computational resources or network services for consumer IoT devices, such as cloud computing, cloud storage service providers, and service providers (e.g., CDN). Third party refers to companies contacted by the device and not categorized as first or support parties. We adopt the methodology from Ren et al. [5] to determine server affiliations through a three-step verification process: (1) We first identify the organization of the IP addresses contacted by the device. If the IP’s organization matches the manufacturer or companion app of the consumer IoT device, we classify it as a first party; (2) For unresolved cases, examine official Privacy Policies and Third-Party Data Sharing documents (prioritizing device-specific policies) to identify support party; (3) Categorize any undocumented or unaffiliated servers as the third party.
@@ 3 Destination Analysis
For consumer IoT devices in China, traffic directed to overseas locations or third-party destinations is of particular concern, as it may increase security and privacy risks compared to traffic flowing to first-party or support-party servers. This section starts with an analysis of traffic destinations, providing a detailed assessment of security and privacy risks through two key aspects: (1) the geographical distribution of traffic destinations by country and (2) the affiliation analysis of contacted server parties.
@@ 3.1 Distribution of Traffic Destinations
This subsection analyzes the country-level distribution of device traffic destinations, focusing on the proportion of traffic to different countries across device types and lifecycle phases.
We first collected IP addresses interacting with devices and queried their geographical affiliations using the IP geolocation API  [12]. Then, we calculate the proportion of traffic going to different countries. In the calculation, we observed that the traffic packet size of cameras, doorbells, and speakers is significantly larger than that of sensors and other devices. If only the proportion of bytes flowing to devices in different countries is counted, there will be a more serious bias. We processed a special calculation method to eliminate the above bias when calculating the traffic proportion.
Let $X$ represent the set of type $A$ device indices in the dataset. For each device $i$ ($i\in X$), let $M_{i}$ denote the traffic volume (in bytes) transmitted to the country $T$, and let $N_{i}$ denote the total traffic volume transmitted to all destinations. The total normalized traffic from type $A$ devices to country $T$ is then given by $A_{T}=\sum_{i\in X}\frac{M_{i}}{N_{i}}$. Further, let $Sum_{A}$ represent the total traffic volume of all type $A$ devices. The proportion of traffic from type $A$ devices directed to country $T$ is computed as: $\frac{A_{T}}{Sum_{A}}$. Similarly, let $Sum_{T}$ represent the total traffic volume from all devices to the country $T$, and $Sum_{all}$ denote the total traffic volume from all devices to all destinations. The overall proportion of traffic directed to the country $T$ across all devices is then given by: $\frac{Sum_{T}}{Sum_{all}}$.
We used Chiplot [34] to generate a Sankey diagram (Figure 1) to visualize the distribution of traffic destinations. In terms of destination, 99.75% of all device traffic remained within China, with only 0.25% directed overseas. From the perspective of device types, hubs, humidifiers, lights, sensors, and plugs did not show overseas traffic, likely due to their simple communication with cloud servers. A total of 32 devices showed overseas traffic, including cameras, speakers, doorbells, TV, mirror, and pet feeder. For each type of device, the cameras had the highest proportion of foreign traffic at 0.6%, followed by speakers at 0.03%, doorbells at 0.01%, and the other devices at 0.02%.
@@
The United States is the most common overseas destination, accounting for 0.24% of total traffic. In particular, dns.google.com is the most frequently accessed foreign domain, communicated by 21 devices. Other domains involved in United States traffic include ntp.com, amazonaws.com, and iotcplatform.com, supporting services such as time synchronization (e.g., 1.cn.pool.ntp.org), remote access based on TUTK (e.g., all-d-master-tutk.iotcplatform.com, used by Xiaobai doorbells and Jooan cameras), and cloud computing (e.g., p2ps2.coolkit.cn, accessed by Sonoff cameras). Traffic to other countries accounted for less than 0.01% of the total and involved 18 domain names. Among them, 10 domains (e.g., ntp.org, time.cloudflare.com) are related to NTP services. Other services included Google Cloud (e.g., bc.googleusercontent.com), IoT platform management (e.g., all-master.iotcplatform.com), IP address lookup (e.g., api.ipify.org), and AWS cloud hosting (e.g., eu-central-1.compute.amazonaws.com).
@@
Figure 2: Overseas traffic destinations of devices grouped by lifecycle phases. Device Lifecycle Phases: SU (setup), ID (idle), IR (interaction), DE (deletion). The destinations with values less than 0.1 are magnified.
To analyze the distribution of traffic across the device lifecycle phases, we visualized the proportional traffic destination during the different lifecycle phases in Figure 2. Since over 99% of the traffic terminates within China, the domestic traffic is omitted from the figure. Therefore, the traffic of plug, light, hub, humidifier, and sensor devices is also not included in the figure. The results show that during the idle phase, devices exhibit a relatively high proportion of overseas traffic, and these traffic primarily flow to NTP services. Apart from the doorbell device, during the setup stage, all the devices show a significant amount of overseas traffic and involve diverse services, such as DNS service, IP address lookup, IoT platform management, and NTP service. This means the device configurations rely on more types of overseas traffic. During both the interaction and deletion phases, the frequency of overseas destinations decreases. This likely suggests that the main functionalities of IoT devices are built on domestic services.
@@
3.2 Service Contacted by Devices
During device operation, frequent traffic communication occurs between devices and cloud service providers. We calculated the number of cloud server affiliations contacted by devices in different lifecycle phases. Table  1 shows the average number of service providers contacted by devices of each type.

### v3

@@ Abstract
In recent years, consumer Internet of Things (IoT) devices have become widely used in daily life. With the popularity of devices, related security and privacy risks arise at the same time as they collect user-related data and transmit it to various service providers. Although China accounts for a larger share of the consumer IoT industry, current analyses on consumer IoT device traffic primarily focus on regions such as Europe, the United States, and Australia. Research on China, however, is currently relatively rare. This study constructs the first large-scale dataset about consumer IoT device traffic in China. Specifically, we propose a fine-grained traffic collection guidance covering the entire lifecycle of consumer IoT devices, gathering traffic from 77 devices spanning 38 brands and 12 device categories. Based on this dataset, we analyze traffic destinations and encryption practices across different device types during the entire lifecycle and compare the findings with the results of other regions. Compared to other regions, our results show that consumer IoT devices in China rely more on domestic services and overall perform better in terms of encryption practices. However, there are still 23/40 devices improperly conducting certificate validation, and 2/77 devices use insecure certificate algorithms. To facilitate future research, we open-source our traffic collection guidance and make our dataset publicly available.
@@
Chinese Consumer IoT, Traffic Dataset, Network Traffic Measurement, Privacy, Security
@@ 1 Introduction
The consumer Internet of Things (IoT) industry has rapidly developed in recent years, with various devices becoming widespread and offering numerous conveniences. According to the report released by China Internet Network Information Center [4] in 2025, as of November 2024, the number of IoT end-users had increased to 2.642 billion. However, the rapid spread of devices has also raised growing concerns regarding security and privacy risks [3]. According to the threat report jointly released by Bitdefender and Netgear in 2024 [5], IoT devices, such as TVs, hubs, and cameras, are subjected to over 2.5 million attack attempts on average every day. As consumer IoT devices usually have the ability to collect and transmit data, the data collected by these devices is often transmitted to the manufacturers of these devices and other service providers. At the same time, data transmission is facing strict legal regulations [1] in China. All data collected and transmitted by these devices is contained within the traffic they generate, therefore, it is necessary to measure consumer IoT traffic from security and privacy perspectives.
@@
Researchers have conducted studies on consumer IoT device traffic in various regions, including Europe [7, 8], the United States [9, 10, 11], and Australia [12, 13], and have identified several security and privacy risks associated with these devices. However, according to IDC [2], China’s consumer IoT industry is projected to reach \$300 billion by 2027. Despite this rapid growth, research on consumer IoT device traffic in China remains limited and lacks a comprehensive analysis. The research in China has not made their data publicly available [6], and only contains the traffic of the camera. This research gap drives us to build a large-scale and comprehensive dataset of consumer IoT traffic in China and conduct a measurement of the dataset.
@@
While constructing the dataset, we introduce methodological improvements compared to existing studies. Previous studies have primarily focused on the device interaction phase [9, 17]. However, we observe differences in traffic between different operational phases of the device, and focusing only on the interaction phase, may lead to incomplete analysis results. Therefore, we propose a fine-grained traffic collection guidance that covers the entire lifecycle of the device to address this limitation. The entire lifecycle refers to all phases during normal device usage, including device setup, user interaction, idle, and deletion [14]. Compared with prior works, the traffic we collected shows greater coverage and can support more fine-grained analysis. Furthermore, we open-source our traffic collection guidance and make our dataset publicly available to facilitate further research.
@@
According to our traffic collection guidance, we collected the first part of device traffic between March 1 and May 30, 2024, covering 50 consumer IoT devices. To further expand the dataset, we conducted two additional rounds of traffic collection involving 27 more devices: the second from November 1 to November 29, 2024, and the third from June 7 to June 19, 2025. During our 4.5 months traffic collection experiment, we compiled a dataset of 108 hours of traffic from 77 consumer IoT devices in China, the total size of the dataset is 3 GB. Compared with the existing datasets, our dataset covers the full lifecycle traffic of the devices and contains the largest number of devices within a single region. Furthermore, For devices capable of firmware updates during the experiment, we also collected traffic after the update to compare the differences pre- and post-update. Based on the traffic collected, we analyzed the traffic destinations, encryption practices, and then compared our findings with those of the United States and the United Kingdom [9, 17].
@@
Our study reveals that, regarding traffic destinations, Chinese devices show a higher reliance on domestic services. The distribution of traffic destinations for Chinese devices is mainly directed to China (99.75%) and relies on domestic organizations. In contrast, devices from the United Kingdom have a significantly higher proportion of overseas traffic, with a greater reliance on organizations from the United States. Regarding encryption practices, devices from different regions all show great encryption practices, although some issues remain. In particular, Chinese devices have less PII exposure in unencrypted traffic and a lower percentage of unencrypted traffic compared to devices in the United States and the United Kingdom. However, 23/40 devices still use improper certificate verification, and 2/77 use insecure certificate algorithms.
@@
  • The first Chinese consumer IoT devices traffic dataset. We constructed a dataset that includes 108 hours of traffic from 77 devices across 12 categories and 38 brands. We have publicly released the dataset [^1] https://github.com/NKUHack4FGroup/Lifecycle-Based-Traffic-Dataset to support future research within the community.
@@
  • Security and privacy measurement of Chinese consumer IoT device traffic. Based on the dataset, we developed scripts to analyze the traffic destinations and encryption practices, and compared the results with those from other regions, revealing the characteristics of Chinese consumer IoT devices and regional differences. We have made the source code available online.
@@ 2 Background
Consumer IoT devices typically refer to IoT devices designed for personal or household use, such as cameras, speakers, and lights. These devices can be bound to and controlled by the companion apps, following similar operational phases (i.e., lifecycle). During these phases, users can control the devices through various methods. Meanwhile, consumer IoT devices communicate frequently with cloud servers during operation to support various device functions. This section introduces the concepts and definitions relevant to the lifecycle, control methods, and cloud servers mentioned above.
@@
The lifecycle of consumer IoT devices. Based on a prior work [14], the lifecycle of consumer IoT devices contains four phases: setup, interaction, idle, and deletion. When users first activate a consumer IoT device, the device will enter the setup phase. During this phase, users need to complete operations such as network configuration and bind the device to the companion app. The interaction phase refers to the phase in which users interact with the device, such as controlling light switches and checking camera status. After setting up the device, we refer to the phase in which the user no longer interacts with the device as the idle phase. Finally, the deletion phase refers to the phase when users unbind the device from the companion app, restoring it to its factory settings.
@@
Control methods of consumer IoT devices. Common control methods of consumer IoT devices include four types: Wide Area Network (WAN) control, Local Area Network (LAN) control, physical control, and multimodal control. WAN control refers to scenarios where the controlling smartphone with the companion app and the consumer IoT device are not on the same local network. In this case, commands between the app and the device must be relayed by a cloud server. For example, when users are away from home during travel, they can monitor the indoor environment in real-time using a camera via a companion app. LAN control occurs when the controlling smartphone and the consumer IoT device are on the same local network. In this case, communications between the device and the app can be established by using local protocols, such as Bluetooth. For example, when users are at home, they can use the companion app to adjust the brightness of smart lights. Physical control involves direct interaction with the device by using the physical buttons. For example, users can press physical switch buttons on the device to reset it or power it off. Consumer IoT devices with various sensors support multimodal control through different methods. Examples include users waking up speakers by using voice commands or interacting with devices via triggers from motion sensors.
@@
Cloud servers of consumer IoT devices. When a consumer IoT device is in operation, the data it collects will be shared with various cloud servers. Following the classification method used in the study by Ren et al. [9], we categorize these servers into three groups: first party, support party, and third party. First party refers to the manufacturer of the consumer IoT device or the developer of its companion app. Support party refers to organizations that provide computational resources or network services for consumer IoT devices, such as cloud computing, cloud storage service providers, and service providers (e.g., CDN). Third party refers to companies contacted by the device and not categorized as first or support parties. We adopt the methodology from Ren et al. [9] to determine the server’s party through a three-step verification process: (1) We first identify the organization of the IP addresses contacted by the device. If the IP’s organization matches the manufacturer or companion app of the consumer IoT device, we classify it as a first party server; (2) For unresolved cases, examine official Privacy Policies and Third-Party Data Sharing documents (prioritizing device-specific policies) to identify the support party server; (3) Categorize any undocumented or unaffiliated servers as the third party server.
3 Traffic collection methodology and measurement scope
This section introduces the traffic collection methodology and measurement scope in our study. The traffic collection methodology includes the device selection criteria, experiment testbed configuration, and traffic collection guidance. The measurement focuses on two main dimensions: traffic destinations and encryption practices. We collected the device traffic between March 1 and June 1, 2024, November 1 to November 29, 2024, and June 7 to June 19, 2025. During the traffic collection experiment, a total of 108 hours of traffic were systematically collected from 77 Chinese consumer IoT devices. We also collected traffic from ten devices after firmware updates to study differences in device behavior before and after firmware updates, generating 26 hours of traffic data covering the pre- and post-firmware updates.
3.1 Experiment testbed
Table 1: Device list.
Category Cameras (29) Doorbells (3) Hubs (9) Humidifiers (2) Lights (6) Plugs (12) Sensors (3) Speakers (9) Other devices (4)
Devices
XiaobaiY2
XiaobaiY3
Xiaomi
Huawei haique
Xiaopai
Cubetoou
ZTE
Yingshi C6CN
360 6C
Jooan
Konka
Pisen
Lenovo
Yibang
Tenda
Mercury
Aqara
Xiaovv
Huawei xiaotun
Xiaotundangjia
TP-Link
IMOU
Greatwall
Xiaomi 3
JPLAYER
Haier
Sonoff
Xiaomao
Xiaobai Smart
IMOU
Chuangmi
TP-LINK
Aqara
Yingshi
Mindor
Tuya
Xiaomi
Jingxun
Meian
Sonoff
BroadLink
Smartmi
Mijia
Huawei dalen
Wiz
Philips
Mijia
Aqara
Midea
Hongyar
Mijia2
Mijia3
Gosund
Yingshi
Mindor
Tuya
Xiaodu
Xiangrikui
Bull
Aqara
Tuya
Yingshi
Xiaodu
Midea Speaker
Huawei Speaker
Xiaomi Speaker
Xiaomi Speaker Plus
Tmall Speaker
Xiaodu Clock
Jingyuzuo Clock
Xiaodu Speaker
XiaomiTV Speaker
Dogness Feeder
Deebot Cleaner
Midea Mirror
Huawei TV
We divide the experiment environment into LAN and WAN to capture device traffic under different network scenarios. To ensure diverse and comprehensive traffic collection, we selected a total of 77 consumer IoT devices across 12 categories (cameras, plugs, speakers, hubs, humidifiers, lights, doorbells, sensors, mirror, TV, Cleaner, and pet feeder). To ensure the representativeness of adopted consumer IoT devices, we prioritized brand recognition, user reviews, and ratings on major Chinese e-commerce platforms (e.g., Taobao, JD.com) during the selection process. The complete list of devices and corresponding brand information is provided in Table 1.
To minimize external interference, experiments are conducted in a controlled environment. The setting of the controlled experiment environment is shown in Figure 1. The router in the figure is a Xiaomi AX9000 running OpenWrt, with Tcpdump installed on both the WAN and LAN interfaces for traffic capture. An Android smartphone, with the latest versions of companion apps downloaded from the corresponding manufacturer’s official websites, is used during the experiments. For LAN operations, both the smartphone and devices are connected to the local area network, while WAN operations require the smartphone to disconnect from the LAN. All devices are configured according to the standard setup process and bind to their respective companion apps. A Windows 10 PC is employed to configure the router and collect the traffic throughout all phases. The traffic collection process follows a semi-automated approach, and the experimenters adhere to the Traffic Collection Guidance (section 3.2). This guidance combines automated scripts with manual operations, ensuring flexibility in the experiment while simulating realistic scenarios.
Figure 1: The setting of the controlled experiment environment.
3.2 Traffic collection guidance
Previous studies have mainly relied on manual methods for traffic collection [15]. While effective, they often lack a standard operating procedure. To address these limitations, Ren et al. [9] proposed a fully automated click-based method to reduce human intervention. However, some operations, like physical controls, remain challenging to automate entirely, and complete automation fails to replicate real-world user behavior accurately. To prevent the effect of personal operations and guarantee comprehensive traffic collection, we propose a semi-automated traffic collection guidance. Our traffic collection guidance follows a semi-automated design principle comprising four core files:
  • Manual operation process files. The file predefined the operations that the experimenters need to perform for each type of device.
  • Timestamp files. Record the operations performed on the device and the corresponding times.
  • Interactive traffic collection script. Start Tcpdump to collect traffic and automatically remind the experimenters to perform corresponding operations according to the manual operation process files. During the collection, the operations and the corresponding times will be recorded in timestamp files. After collecting the traffic, the raw traffic will be saved.
  • Traffic segmentation script. Split the raw traffic according to Timestamp files.
Upon initiating traffic capture, the collection script starts Tcpdump and loads the manual operation process file. Experimenters are guided by operation prompts to perform predefined operations, such as device setup and firmware updates. The start and end times of each operation are automatically logged in the timestamp file. After traffic collection, the segmentation script processes the raw PCAP file using the timestamp file, automatically splitting it into multiple lifecycle phases for analysis. Additionally, there are 13% of all devices that supported firmware updates during the experiment, including five cameras (Aqara camera, TP-Link camera, XiaobaiY2 camera, Xiaomi camera), four hubs (Aqara hub, Boardlink hub, Xiaomi hub, Yingshi hub), and one plug device (Mijia plug). For these devices, we collected traffic both before and after the update process. After the update, devices are re-bound to their companion apps, and traffic is recaptured using the same process. This semi-automated approach ensures comprehensive coverage of the device lifecycle, accurately reflecting real-world controls. Following the established environment and traffic collection guidance, we successfully collected 108 hours of traffic, including 13 hours of post-firmware-update traffic from 10 devices.
3.3 Measurement scope
The consumer IoT devices continuously collect large amounts of data about users and their surrounding environments during daily use. Much of this data may contain sensitive personal information, making it crucial to evaluate how these devices transmit data and to what destinations. Our study focuses on the security and privacy practices of these devices by analyzing both the traffic destinations and encryption practices.
First, we analyzed the traffic destinations of the devices. Given the location of the devices, all the devices analyzed in this study are popular products in China, therefore, we focus on traffic directed to destinations outside of China. Through traffic analysis, we assess where devices send data, the proportion of domestic versus overseas traffic, and the services provided by these destinations. Additionally, we further analyze which parties the service providers belong to because the traffic that goes to non-first parties tends to have a higher risk than traffic that goes to first parties. Moreover, we identified several domains frequently accessed by most devices, which were linked to specific organizations. By analyzing these domains, we identified the organizations most frequently contacted by the devices. This analysis sheds light on the reliance of consumer IoT devices on different organizations.
Second, we conduct a detailed analysis of the encryption practices of device traffic. First, we measure the proportion of encrypted traffic. We then categorize the types of encryption protocols used in the encrypted traffic. Due to TLS 1.2 is widely used in device traffic, we conduct an in-depth analysis of the security of certificates using TLS 1.2, focusing on whether the algorithms used in the certificates meet current security standards and whether the device verifies the certificate correctly. For unencrypted traffic, we assess the potential risk of exposing personally identifiable information (PII) in plain text, with particular attention to the potential leakage of sensitive information.
@@ 4 Destination analysis
For consumer IoT devices in China, traffic directed to overseas locations warrants attention. Similarly, traffic flows to third party servers generally present higher security and privacy risks compared to traffic flows to first party or support party servers. This section starts with an analysis of traffic destinations, providing a detailed assessment of security and privacy risks through two key aspects: (1) the geographical distribution of traffic destinations by country and (2) the analysis of contacted server parties.
@@ 4.1 Distribution of traffic destinations
This subsection analyzes the country-level distribution of device traffic destinations, focusing on the proportion of traffic to different countries in different device types and lifecycle phases.
We first collected IP addresses interacting with devices and queried their geographical locations using the IP geolocation API  [16]. Then, we calculate the proportion of traffic going to different countries. In the calculation, we observed that the traffic packet size of cameras, doorbells, and speakers is significantly larger than that of sensors and other devices. If only the proportion of bytes flowing to devices in different countries is counted, the final result will largely depend on the traffic destinations of devices such as cameras, and the influence of sensors and other devices will be negligible, resulting in a bias. To eliminate the above bias, we processed a new calculation method.
Let $X$ represent the set of type $A$ device indices in the dataset. For each device $i$ ($i\in X$), let $M_{i}$ denote the traffic (in bytes) transmitted to the country $T$, and let $N_{i}$ denote the total traffic transmitted to all destinations. The total normalized traffic from type $A$ devices to country $T$ is then given by $A_{T}=\sum_{i\in X}\frac{M_{i}}{N_{i}}$. Further, let $Sum_{A}$ represent the total traffic of all type $A$ devices. The proportion of traffic from type $A$ devices directed to country $T$ is computed as: $\frac{A_{T}}{Sum_{A}}$. Similarly, let $Sum_{T}$ represent the total traffic from all devices to the country $T$, and $Sum_{all}$ denote the total traffic from all devices to all destinations. The overall proportion of traffic directed to the country $T$ across all devices is then given by: $\frac{Sum_{T}}{Sum_{all}}$.
We used Chiplot [38] to generate a Sankey diagram (Figure 2) to visualize the distribution of traffic destinations. In terms of destination, 99.75% of all device traffic remained within China, with only 0.25% directed overseas. From the perspective of device types, hubs, humidifiers, lights, sensors, and plugs did not show overseas traffic, likely due to their simple communication with cloud servers. A total of 32 devices showed overseas traffic, including cameras, speakers, doorbells, TV, mirror, and pet feeder. For each type of device, the cameras had the highest proportion of overseas traffic at 0.6%, followed by speakers at 0.03%, doorbells at 0.01%, and the other devices at 0.02%.
@@
The United States is the most common overseas destination, accounting for 0.24% of total traffic. In particular, dns.google.com is the most frequently accessed overseas domain, communicated by 21 devices. Other domains involved in the United States traffic include ntp.com, amazonaws.com, and iotcplatform.com, supporting services such as time synchronization (e.g., 1.cn.pool.ntp.org), remote access based on TUTK (e.g., all-d-master-tutk.iotcplatform.com, used by Xiaobai doorbells and Jooan cameras), and cloud computing (e.g., p2ps2.coolkit.cn, accessed by Sonoff cameras). Traffic to other countries accounted for less than 0.01% of the total and involved 18 domain names. Among them, 10 domains (e.g., ntp.org, time.cloudflare.com) are related to NTP services. Other services included Google Cloud (e.g., bc.googleusercontent.com), IoT platform management (e.g., all-master.iotcplatform.com), IP address lookup (e.g., api.ipify.org), and AWS cloud hosting (e.g., eu-central-1.compute.amazonaws.com).
@@
Figure 3: Overseas traffic destinations of devices grouped by lifecycle phases. Device lifecycle phases: SU (setup), ID (idle), IR (interaction), DE (deletion). The destinations with values less than 0.1 are magnified. The vertical coordinate value is the ratio of the number of bytes flowing from the device to a certain country to the total number of bytes of the device’s flow, multiplied by the number of devices.
To analyze the distribution of traffic across the device lifecycle phases, we visualized the proportional traffic destination during the different lifecycle phases in Figure 3. Since over 99% of the traffic terminates within China, the domestic traffic is omitted from the figure. Therefore, the traffic of plug, light, hub, humidifier, and sensor devices is also not included in the figure. The results show that during the idle phase, devices exhibit a relatively high proportion of overseas traffic, and the traffic primarily flows to NTP services. Apart from the doorbell device, during the setup phase, all the devices show a significant amount of overseas traffic and involve diverse services, such as DNS service, IP address lookup, IoT platform management, and NTP service. This means the device configurations rely on many types of overseas services. During both the interaction and deletion phases, the frequency of overseas destinations decreases. This likely suggests that the main functionalities of IoT devices are built on domestic services.
@@
4.2 Servers contacted by devices
During device operation, frequent traffic communication occurs between devices and cloud service providers. We calculated the number of cloud servers belonging to different parties contacted by devices in different lifecycle phases. Table 2 shows the average number of service providers contacted by devices of each type.